n=1.35 n=135 to a medium with an indax of retra cton o4 15. T the incident ande is 30 what is the refrac η_1.5 ted angle? A ray of kght gots froma medior rth index of refraction of 13 to an unknovon index of refradio Use the anqlus in the diagrar to find the untnousn nz 20 /n-1:3
Answer #1

1) using snell's law

1.35sin30 = 1.5sin\theta

=>  \theta = sin-1(0.45) = 26.7437o

2) using snell's law,

1.3sin(\theta) = n2sin20

=> n2 = 3.8009sin(\theta)

Note:  \theta is not visible in the picture, you can get n2 by just plug in the \theta value in the above relation.